# Linear Attention for Joint Power Optimization and User-Centric Clustering in Cell-Free Networks

### v1

@@ V-B Training approach
Unlike standard supervised learning [nasteski2017overview], the optimal powers cannot be precomputed before training because the clusters of APs serving each UE are determined dynamically by the model itself during each training epoch. The overall training procedure is summarized in Algorithm 1, which illustrates the dynamic supervision process where labels are generated on‑the‑fly based on intermediate model decisions.
During the forward pass of a training epoch, the model evaluates attention scores between the input features. These scores define the AP clusters, assigning to each UE a set of serving APs. Once the clusters are determined, the UL and DL power control problems are solved for the predicted clusters using the closed‑form expression in [miretti2022closed]. The resulting optimal powers are then used in the backward pass to guide training. Specifically, the predicted powers $\tilde{\mathbf{p}}$ are compared with the normalized optimal ones $\tilde{\mathbf{p}}^{\star}$, and the loss function updates the model parameters:
$\!\!\!\!\text{Loss}=\frac{1}{S}\sum_{i=1}^{S}\|\tilde{\mathbf{p}}^{\star}_{i}\!-\!\tilde{\mathbf{p}}_{i}\|^{2}-\lambda\left(\min_{k}\textrm{\acs{SE}}_{k}^{\text{UL}}+\min_{k}\textrm{\acs{SE}}_{k}^{\text{DL}}\right).\!\!$ (32)
The loss combines the mean‑squared error between predicted and optimal powers with a clustering‑aware penalty term that promotes fairness across users. The hyperparameter $\lambda$ controls the trade‑off: larger values emphasize balancing the worst‑case SE, while smaller values prioritize minimizing prediction error. In practice, $\lambda$ is tuned empirically via cross‑validation to achieve a balance between accuracy and fairness.
The model is implemented in PyTorch [paszke2019pytorch] and configured with the parameters listed in Tables I and II, empirically tuned to balance accuracy and computational efficiency.
Our training strategy follows a supervised learning paradigm with online label generation, also referred to as learning with dynamic supervision [pham2021meta, wang2020learning]. In this framework, ground-truth labels are computed on-the-fly during training based on the model’s intermediate decisions. Notably, the model does not take channel coefficients as input; instead, their effect is encapsulated in the optimal power labels. By learning directly from these labels, the network implicitly captures the propagation characteristics of the environment and learns to map UE-AP spatial configurations to clustering and power-allocation decisions. As a result, the model effectively approximates the max-min SE solution at inference time without requiring any explicit optimization.
@@
Trade‑off parameter $\lambda$ $10^{-2}$
@@
Optimizer AdamW [zhou2024towards]
@@
In this section, we assess the performance of the proposed learning-based framework for solving the MMF problem in the user-centric cell-free mMIMO system described in Sect. II. The main simulation parameters are reported in Table III. All numerical results are averaged over the test set.
The proposed solution is compared against the following clustering and power-allocation schemes:
  • Optimal CF: A fully cooperative baseline in which all APs serve all UEs, and transmit powers are obtained via the closed-form expression in [miretti2022closed].
  • DCC: A fixed-clustering benchmark based on the dynamic cooperation clustering (DCC) scheme [bjornson2019new], where each UE is associated with its $Q=8$ strongest APs. Power control follows the closed-form solution.
@@
Table III: Cell-free mMIMO network parameters.
@@
Path‑loss exponent $3.67$
@@
Shadow fading $F_{kl}$ $\mathcal{N}_{\mathbb{C}}(0,\alpha^{2}),\;\alpha^{2}=4\,\mathrm{dB}$
@@
VI-A Spectral efficiency evaluation
@@
Figure 6 reports the average per-UE SE as a function of the number of users $K$ for both UL and DL transmission. As expected, the per-UE SE decreases with increasing $K$ in both links, since a larger number of users must share a fixed amount of AP resources and is subject to higher interference. Nevertheless, the rate at which performance degrades varies significantly across the considered strategies. In the UL (Figure 6(a)), Optimal CF provides the highest SE for small values of $K$ due to full AP cooperation. However, its performance deteriorates more rapidly as $K$ increases. Beyond $K\approx 10$, Optimal UC-CF surpasses the fully cooperative baseline, as limiting cooperation to the most relevant APs reduces interference and enables a more effective allocation of transmit power. A similar behavior is observed in the DL (Figure 6(b)): while full cell-free (CF) remains superior at low user densities, Optimal UC-CF achieves higher SE for larger $K$, with the benefits of clustering being even more pronounced in DL due to improved power concentration and reduced inter-user interference. The Predicted UC-CF closely tracks the optimal user-centric (UC)-CF in both UL and DL, confirming the generalization capability of the trained model across unseen user loads. In contrast, the classical DCC consistently yields the lowest SE, as its fixed clustering cannot adapt to varying interference conditions.
Overall, these results show that while fully cell-free operation is advantageous at low UE densities, user-centric clustering becomes increasingly beneficial as the network grows congested. The proposed learning model adapts to this transition, sustaining near-optimal performance with reduced computational and signaling overhead.
VI-B AP–UE clustering performance
@@
VI-C Power prediction accuracy
To further evaluate the effectiveness of the proposed learning model, we compare the cumulative distribution functions of UL and DL powers obtained from the optimal closed-form solution [miretti2022closed] and those predicted by the trained model on the test set. As shown in Figure 9, the resulting curves are nearly overlapping in both cases, indicating that the model successfully learns the underlying power allocation strategy for the predicted clusters. This close agreement demonstrates that the predicted powers not only approximate the optimal allocation with high fidelity but also preserve the statistical distribution of transmit powers across UEs. Consequently, the proposed approach achieves reliable power prediction while avoiding the computational burden of solving the closed-form optimization problem.
@@ VI-C Power prediction accuracy
VI-D Robustness to input errors
@@
In Figure 10, we analyze the robustness of the proposed learning model against input uncertainty. During training, the model is exposed to user position inputs corrupted by random Gaussian noise with a standard deviation of $1\,\mathrm{m}$. To assess generalization and robustness, the model is evaluated under three test conditions: i) input with a fixed standard deviation error of $3\,\mathrm{m}$, ii) input with a fixed standard deviation error of $5\,\mathrm{m}$, and iii) an ideal scenario with error-free input.
@@
VI-E Comparison with other Transformer variants
@@
Figure 11 presents a comparative analysis of UL and DL per-UE SE achieved by the Transformer and four linear attention variants. Linformer is excluded because it cannot accommodate variable UE counts without additional preprocessing or architectural modifications [wang2020linformer].
In both UL and DL scenarios, ELU-CosFormer consistently outperforms the other models across all user counts. This confirms the effectiveness of integrating the ELU activation into the CosFormer kernel. CosFormer also shows strong performance, closely trailing ELU-CosFormer, while Performer and Transformer achieve moderate SE levels. Longformer, although competitive at low user counts, exhibits a steeper decline as $K$ increases, indicating reduced scalability. The superior performance of ELU-CosFormer stems from the integration of ELU with CosFormer’s cosine-based attention kernel, as explained in Section IV. As a result, ELU-CosFormer forms physically consistent clusters and outperforms other variants across all user counts under our fairness-driven objective.
@@
  • Memory (MB): Estimated memory footprint during inference, reflecting GPU usage.
@@
ELU-CosFormer and CosFormer share the same architecture and parameter count, differing only in their activation functions-ELU and ReLU, respectively. Although ELU is computationally more intensive, ELU-CosFormer achieves slightly lower training time per epoch, likely due to smoother gradient flow that enables more efficient backpropagation and faster convergence. In contrast, ReLU’s simplicity gives CosFormer a marginal advantage in inference latency, as ReLU operations are highly optimized for GPU execution. Performer maintains a low parameter count and memory footprint comparable to CosFormer, but exhibits higher training and inference times due to the computational overhead of random feature projections and exponentiation in its kernel mapping [choromanski2020rethinking], which are less GPU-efficient than CosFormer’s cosine reweighting.
@@
When the number of users $K$ increases, the per-sample computational complexity of our model is dominated by the linearized attention term $\mathcal{O}(M\cdot d_{\text{mod}}\cdot K)$, which scales linearly with $K$. This linear scaling represents a clear advantage compared to prior Transformer-based approaches [kocharlakota2024pilot, chafaa2025transformer], where the complexity grows quadratically as $\mathcal{O}(M\cdot d_{\text{mod}}\cdot K^{2})$, and iterative optimization methods [farooq2020accelerated], which scale as $\mathcal{O}(T\cdot L\cdot K^{2})$ with $T$ iterations. It is also favorable compared to the optimal closed-form solution that require cubic scaling $\mathcal{O}(K^{3})$, which quickly become prohibitive in dense networks.
@@
Beyond asymptotic complexity, several practical aspects reinforce the advantage of our model. The linear attention mechanism reduces memory usage compared to quadratic approaches, enabling scalability to larger user counts and deployment on resource‑constrained edge devices. This efficiency also translates into lower energy consumption, aligning with sustainability goals in next‑generation wireless networks. Moreover, the one‑shot inference design avoids iterative overhead, ensuring real‑time responsiveness to dynamic user distributions and mobility. Importantly, the model maintains near‑optimal SE despite its reduced complexity, confirming that efficiency does not compromise accuracy. Finally, linear attention benefits more from parallelization on graphics processing unit and tensor processing unit [pal2019gpuparallel, jouppi2017tpu], further widening the performance gap with quadratic and iterative methods.
@@
This paper proposed a scalable and flexible deep learning model for joint AP clustering and power allocation in user-centric cell-free massive MIMO systems. By relying only on spatial information and leveraging a customized CosFormer attention mechanism, our approach eliminates the need for channel estimation, avoids pilot contamination, and adapts seamlessly to dynamic network configurations. Numerical evaluations demonstrate that the proposed method achieves near-optimal SE with significantly reduced computation and signaling overhead, making it a practical solution for dynamic and large-scale systems.
Future research directions include investigating distributed learning strategies among APs or clusters to further reduce centralization requirements. Another promising direction is the integration of temporal information, enabling the model to capture user mobility, traffic variations, and time-dependent channel dynamics. Extending the architecture with temporal attention mechanisms may enable proactive resource allocation and enhance robustness in highly dynamic operating conditions.

### v2

@@ V-B Training approach
Unlike conventional supervised learning [25], the optimal power labels cannot be precomputed prior to training because the clusters of APs serving each UE are determined dynamically by the model itself. Consequently, the optimal power labels depend on the clustering decisions produced during the forward pass and are computed on-the-fly during training. This results in a supervised learning framework with dynamic label generation, also referred to as dynamic supervision [31, 37], where the supervision signals are obtained by solving the power control problems for the clusters predicted by the model.
The overall training procedure is summarized in Algorithm 1. During the forward pass, the model evaluates attention scores between the input features, which define the AP clusters by assigning a set of serving APs to each UE. Given these predicted clusters, the UL and DL power control problems are solved using the closed-form expressions in [24]. Then, the predicted powers $\tilde{\mathbf{p}}$ are compared with the normalized optimal ones $\tilde{\mathbf{p}}^{\star}$, and the loss function updates the model parameters:
$\!\!\!\!\mathcal{L}=\frac{1}{S}\sum_{i=1}^{S}\|\tilde{\mathbf{p}}^{\star}_{i}-\tilde{\mathbf{p}}_{i}\|^{2}-\lambda\left(\min_{k}\textrm{\acs{SE}}_{k}^{\text{UL}}+\min_{k}\textrm{\acs{SE}}_{k}^{\text{DL}}\right),$ (30)
where $S$ is the batch size. The loss combines the mean-squared error between predicted and optimal powers with a clustering-aware penalty that promotes fairness across users. The hyperparameter $\lambda$ controls the trade-off between minimizing prediction error and improving the worst-case SE. In practice, $\lambda$ is tuned empirically via cross-validation.
The model is implemented in PyTorch [16] and configured with the parameters listed in Tables I and II, which are empirically tuned to balance accuracy and computational efficiency.
Notably, the model does not take channel coefficients as input; instead, their effect is captured through the optimal power labels. By learning from these labels, the network maps UE-AP spatial configurations to clustering and power allocation decisions across different network configurations and user loads, effectively approximating the max-min SE policy at inference time without requiring instantaneous channel information.
@@
Trade-off parameter $\lambda$ $10^{-2}$
@@
Optimizer AdamW [39]
@@
In this section, we assess the performance of the proposed learning-based framework for solving the MMF problem in the UC CF mMIMO system described in Sect. II. The main simulation parameters are reported in Table III. All numerical results are averaged over the test set.
The proposed solution is compared against the following clustering and power allocation schemes:
  • Optimal CF: A fully cooperative baseline in which all APs serve all UEs, and transmit powers are obtained via the closed-form expression in [24].
  • DCC: A fixed-clustering benchmark based on the DCC scheme [6], where each UE is associated with its $Q=8$ strongest APs. Power control follows the closed-form solution.
@@
[…]
@@
Path-loss exponent $3.67$
@@
Shadow fading $F_{kl}$ $\mathcal{N}_{C}(0,\alpha^{2}),\;\alpha^{2}=4\,\mathrm{dB}$
@@
VI-A Activation function performance
To validate the choice of the ELU activation function detailed in Sect. IV-B, we evaluate the per-UE SE achieved by a CosFormer using several activation functions [33] for the attention computation with the simulation parameters in Table III. The results reported in Figure 4 demonstrate that ELU consistently yields higher SE across all user counts $K$ for both UL and DL scenarios, confirming that ELU provides a more effective attention computation for our problem.
(a) UL
(b) DL
VI-B Spectral efficiency evaluation
@@
Figure 5 reports the average per-UE SE as a function of the number of users $K$ for both UL and DL transmission. As expected, the per-UE SE decreases with increasing $K$ in both links, since a larger number of users must share a fixed amount of AP resources and is subject to higher interference. Nevertheless, the rate at which performance degrades varies significantly across the considered strategies. In the UL (Figure 5(a)), Optimal CF provides the highest SE for small values of $K$ due to full AP cooperation. However, its performance deteriorates more rapidly as $K$ increases. Beyond $K\approx 10$, Optimal UC-CF surpasses the fully cooperative baseline, as limiting cooperation to the most relevant APs reduces interference and enables a more effective allocation of transmit power. A similar behavior is observed in the DL (Figure 5(b)): while full CF remains superior at low user densities, Optimal UC-CF achieves higher SE for larger $K$, with the benefits of clustering being even more pronounced in DL due to improved power concentration and reduced inter-user interference. The Predicted UC-CF closely tracks the optimal UC-CF in both UL and DL, confirming the generalization capability of the trained model across unseen user loads. In contrast, the classical DCC consistently yields the lowest SE, as its fixed clustering cannot adapt to varying interference conditions.
Overall, these results show that while fully CF operation is advantageous at low UE densities, UC clustering becomes increasingly beneficial as the network grows congested. The proposed learning model adapts to this transition, sustaining near-optimal performance with reduced computational and signaling overhead.
VI-C AP-UE clustering performance
@@
VI-D Power prediction accuracy
To further evaluate the effectiveness of the proposed learning model, we compare the cumulative distribution functions of UL and DL powers obtained from the optimal closed-form solution [24] and those predicted by the trained model on the test set. As shown in Figure 8, the resulting curves are nearly overlapping in both cases, indicating that the model successfully learns the underlying power allocation strategy for the predicted clusters. This close agreement demonstrates that the predicted powers not only approximate the optimal allocation with high fidelity but also preserve the statistical distribution of transmit powers across UEs. Consequently, the proposed approach achieves reliable power prediction while avoiding the computational burden of solving the closed-form optimization problem. This behavior is explained by the fact that the optimal power allocation is a mapping of large-scale channel statistics which highly depend on UE-AP geometry, and thus the trained model learns to approximate this mapping accurately even for unseen UE deployments of the test set.
@@ VI-D Power prediction accuracy
VI-E Robustness to input errors
@@
In Figure 9, we analyze the robustness of the proposed learning model against input uncertainty. During training, as stated in (28), the model is exposed to user position inputs corrupted by random Gaussian noise with a standard deviation of $1\,\mathrm{m}$. To assess generalization and robustness, the model is evaluated under three test conditions: i) input with a fixed standard deviation error of $3\,\mathrm{m}$, ii) input with a fixed standard deviation error of $5\,\mathrm{m}$, and iii) an ideal scenario with error-free input.
@@
VI-F Comparison with other Transformer variants
@@
Figure 10 presents a comparative analysis of UL and DL per-UE SE achieved by the Transformer and four linear attention variants. Linformer is excluded because it cannot accommodate variable UE counts without additional preprocessing or architectural modifications [36].
In both UL and DL scenarios, ELU-CosFormer consistently achieves the highest SE across all user counts. This result highlights the benefit of integrating the ELU activation into the CosFormer attention kernel. As discussed in Sect. IV, the combination of ELU with the cosine-based attention mechanism improves the model’s ability to capture the spatial interactions between UEs and APs. Consequently, ELU-CosFormer produces more effective AP clustering and power allocation decisions, maintaining strong performance even as the number of users increases under the considered fairness-driven objective.
@@
  • Memory (MB): Estimated memory footprint during inference, reflecting graphics processing unit (GPU) usage.
@@
ELU-CosFormer and CosFormer share the same architecture and parameter count, differing only in their activation functions-ELU and ReLU, respectively. Although ELU is computationally more intensive, ELU-CosFormer achieves slightly lower training time per epoch, due to smoother gradient flow that enables more efficient backpropagation and faster convergence. In contrast, ReLU’s simplicity gives CosFormer a marginal advantage in inference latency, as ReLU operations are highly optimized for GPU execution. Performer maintains a low parameter count and memory footprint comparable to CosFormer, but exhibits higher training and inference times due to the computational overhead of random feature projections and exponentiation in its kernel mapping [12], which are less GPU-efficient than CosFormer’s cosine reweighting.
@@
When the number of users $K$ increases, the per-sample computational complexity of our model is dominated by the linearized attention term $\mathcal{O}(M\cdot d_{\text{mod}}\cdot K)$, which scales linearly with $K$. This linear scaling represents a clear advantage compared to prior Transformer-based approaches [20, 10], where the complexity grows quadratically as $\mathcal{O}(M\cdot d_{\text{mod}}\cdot K^{2})$, and iterative optimization methods [17], which scale as $\mathcal{O}(T\cdot L\cdot K^{2})$ with $T$ iterations. It is also favorable compared to the optimal closed-form solution that require cubic scaling $\mathcal{O}(K^{3})$, which quickly become prohibitive in dense networks.
@@
Beyond asymptotic complexity, several practical aspects reinforce the advantage of our model. The linear attention mechanism reduces memory usage compared to quadratic approaches, enabling scalability to larger user counts and deployment on resource-constrained edge devices. This efficiency also translates into lower energy consumption, aligning with sustainability goals in next-generation wireless networks. Moreover, the one-shot inference design avoids iterative overhead, ensuring real-time responsiveness to dynamic user distributions and mobility. Importantly, the model maintains near-optimal SE despite its reduced complexity, confirming that efficiency does not compromise accuracy. Finally, linear attention benefits more from parallelization on graphics processing unit and tensor processing unit [30, 18], further widening the performance gap with quadratic and iterative methods.
@@
This paper proposed a scalable and flexible deep learning model for joint AP clustering and power allocation in UC CF mMIMO systems. The method follows a dynamic supervised learning paradigm in which optimal power labels are generated on the fly for the predicted AP clusters to train a CosFormer-based model that maps UE-AP spatial configurations to the corresponding resource allocation decisions. By relying primarily on spatial information and leveraging a customized attention mechanism, the proposed framework avoids pilot contamination, adapts seamlessly to varying user loads, and enables resource allocation decisions without requiring instantaneous channel information as input. This allows the joint clustering and power control to be decoupled from the data detection stage.
Numerical evaluations demonstrated that the proposed framework achieves near-optimal SE while reducing computational complexity and signaling overhead. In particular, while the complexity of the optimal solution scales as $\mathcal{O}(K^{3})$ with the number of UEs, the proposed model performs resource allocation through a single inference step and scales linearly with $K$. Furthermore, since the model relies only on spatial information during inference, it eliminates the need for exchanging instantaneous channels for resource optimization, thereby reducing signaling overhead. In addition, the clustering learned by the model limits the number of APs serving each UE to the most relevant ones, which further reduces coordination and signaling requirements between APs and the CPU.
Future research directions include investigating distributed learning strategies among APs or clusters to further reduce centralization requirements. Another promising direction is the integration of temporal information, enabling the model to capture user mobility, traffic variations, and time-dependent channel dynamics. Extending the architecture with temporal attention mechanisms may enable proactive resource allocation and enhance robustness in highly dynamic operating conditions. Additionally, further validation could use ray-tracing datasets or real measurement campaigns.